# GAF: Gaussian Avatar Reconstruction from Monocular Videos via Multi-view Diffusion

### v1

@@ 5.2 Ablation Studies
(a) (b) (c) (d) (e) (f) (g) (h) (i)
Figure 6: Ablation Studies on different types of diffusion priors. (a) Input; (b) Ground truth; Comparisons between method variants of (c) No diffusion; using (d) Pretrained Stable Diffusion; (e) Personalized Stable Diffusion; (f) Pose-conditioned multi-view diffusion; (g) Our multi-view diffusion using Score Distillation Sampling (SDS) loss; (h) Ours without latent upsampler $\times$2; (i) Ours. Our normal map-conditioned multi-view diffusion priors enable more photo-realistic novel views with identity and appearance consistency, by constraining novel views using pseudo-image ground truths, which are decoded from iteratively denoised latents followed by a latent upsampler.
@@
We conduct detailed ablation studies to verify the effectiveness of each design in our multi-view head diffusion prior learning. We select six sequences from NeRSemble dataset, including ’055 EXP-5’, ’098 EMO-1’, ’134 EMO-1’, ’165 EMO-1’, ’221 EXP-8’, and ’417 EMO-4’. The results are presented in Figure 6, and Table. 3.
@@
An alternative is to use pre-trained text-to-image diffusion models Stable Diffusion [52]. Another alternative is to fine-tune Stable Diffusion on RGB frames from the input video, obtaining personalized image diffusion priors. We implement both variants using normal map guidance of ControlNet [84]. In Fig. 6 (d) , pre-trained text-to-image priors often produce renderings that deviate from the original identity. In Fig. 6 (e), personalized diffusion priors improve identity preservation but struggle with appearance consistency, as they lack input image information to hallucinate novel views. Our approach learns to jointly generate multiple novel views conditioned on the input image, thus achieving higher realism and view consistency in both identity and appearance.
What is the effect of normal map condition for multi-view diffusion models?
Our multi-view head diffusion models condition on normal maps of FLAME reconstruction, which are pixel-aligned with the target novel-view images. From Fig. 6 (f) reflects that using camera pose conditioning could introduce obvious misalignment errors in synthesizing pseudo-image ground truths, leading to blurred 3D Gaussian renderings.
@@
We can instead use Score Distillation Sampling (SDS) loss to constrain multi-view renderings. As shown in Fig. 6 (g), SDS loss has appearance oversaturation issues in the face region.
@@
Through Fig. 6 (h) vs. (i), we can see that the upsampler module significantly sharpens and enhance facial appearance details.
5.3 Robustness Analysis
@@ 6 Conclusion
In this work, we present a novel method to reconstruct photo-realistic head avatars from monocular videos to push the frontier of avatar fidelity from commodity devices. Due to limited observation and coverage of human heads, Gaussian reconstruction from monocular videos is inherently under-constrained. To address this challenge, we introduce multi-view diffusion priors that jointly constrain photorealism across multiple views rendered from Gaussian splats, while preserving face identity and appearance consistency. We obtain these priors by designing a multi-view head diffusion model, fine-tuned on a multi-view head video dataset to generate novel views from a single image, conditioned on rendered normal maps from FLAME head reconstruction. To prevent appearance over-saturation, we apply an effective diffusion loss using iteratively denoised images as pseudo-ground truths. For finer facial details, we combine a latent upsampler diffusion model with our multi-view diffusion. By reducing data capture requirements for avatar creation, our approach has the potential to unlock new opportunities in immersive VR/AR applications and products.
@@ Acknowledgements
This work was supported by Toyota Motor Europe and Woven by Toyota. This work was also supported by the ERC Starting Grant Scan2CAD (804724). We thank Rundi Wu, Simon Giebenhain, and Lei Li for constructive discussions and Angela Dai for the video voice over.
@@ Appendix A Dataset
The statistics of the monocular video data are summarized in  Tab. 5. Since monocular videos captured on commodity devices lack corresponding ground truths for novel view renderings, we only evaluate avatar animation performance in the quantitative comparisons, by applying pose and expression parameters from unseen frames during training.
@@ B.1 Monocular Head Tracking
The tracking is performed by minizming a combination of multiple energy terms: (i) a photometric energy term, computed between the rendered image and the ground-truth one; (ii) a landmarks energy term, which computes the distance between the projected 2D FLAME [30] landmarks and the 2D landmarks predicted by an off-the-shelf detector [89]; (iii) temporal energy terms, applied on the per-timestep parameters, which ensure smoothness over time; (iv) regularization energy terms, applied on all FLAME [30] parameters. We revised the loss weights for the smoothness terms as: $\lambda_{smooth,transl}=3\mathrm{e}{4}$, $\lambda_{smooth,rot}=3\mathrm{e}{3}$, $\lambda_{smooth,jaw}=4.0$, $\lambda_{smooth,eyes}=1.0$, $\lambda_{smooth,expr}=0.5$. For all the remaining hyper-parameters we refer to the original work [47].
@@ B.2 Multi-view Latent Head Diffusion
Appendix C Generation Results of Multi-view Head Diffusion
@@
Figure 9: Generation sample results of the multi-view head latent diffusion model. Given a single image as input, our method can generate identity-preserved, and view-consistent multi-view portrait images.
@@ Appendix D Additional Comparisons
D.1 Dynamic Head Avatar Reconstruction
In Fig. 11 and Fig. 12, we provide additional qualitative comparisons on dynamic head avatar reconstruction from monocular videos sampled from NerSemble dataset [27].
Figure 10: Network architecture of our multi-view head latent diffusion model. The denoiser network is built on a 2D U-Net architecture with attention blocks. The input consists of multi-view image latents concatenated with VAE latents of normal maps rendered from the FLAME mesh. The 3D Attention block enforce 3D consistency by applying cross-attention across all views. It also incorporates the input image latent into the denoising process, effectively preserving the identity and appearance details of the input portrait.
(a) Input frame (b) Ground truth (c) INSTA (d) FlashAvatar (e) GaussianAvatars (f) Ours
Figure 11: Additional results on novel view synthesis from monocular videos from the NeRSemble dataset. Our method demonstrates robust reconstruction of less observed regions, such as side facial areas, and consistently produces more plausible and view-consistent renderings from hold-out views.
(a) Input (b) Ground truth (c) INSTA (d) FlashAvatar (e) GaussianAvatars (f) Ours
Figure 12: Additional results of head avatar reconstruction from monocular videos on commodity devices. We present novel expression animation results using unseen frames from the monocular videos during training. Additionally, we display the fitting results for the input frame (top right) and novel view renderings of the same frame (bottom right). While all methods accurately fit observed frames in front-facing sequences with limited head poses, baseline methods fail to generalize to novel views and poses due to the absence of effective priors for less unobserved regions.
@@ D.1 Dynamic Head Avatar Reconstruction
D.2 Self- & Cross-Reenactment

### v2

@@ 5.2 Ablation Studies
Input GT (a) (b) (c) (d) (e) (f) (g) (h)
Figure 6: Ablation studies on different types of diffusion priors. Comparisons between method variants of (a) No diffusion; using (b) Pretrained Stable Diffusion; (c) Personalized Stable Diffusion; (d) Pose-conditioned multi-view diffusion; (e) Raymap-conditioned multi-view diffusion; (f) Our multi-view diffusion using Score Distillation Sampling (SDS) loss; (g) Ours without latent upsampler $\times$2; (h) Our final model. Our normal map-conditioned multi-view diffusion priors enable more photo-realistic novel views with identity and appearance consistency, by constraining novel views using pseudo-image ground truths, which are decoded from iteratively denoised latents followed by the latent upsampler.
We conduct detailed ablation studies to evaluate diffusion prior choices for novel view constraints, confirming the effectiveness of each design in our face-specific multi-view diffusion prior learning for 3D head reconstruction from monocular videos. We use six sequences from the NeRSemble dataset. The results are presented in Fig. 6, and Tab. 3. Please refer to the supplementary material for more implementation details of ablation studies.
@@
An alternative is to use pre-trained text-to-image diffusion models Stable Diffusion [64]. Another alternative is to fine-tune Stable Diffusion on RGB frames from the input video, obtaining personalized image diffusion priors. We implement both variants using normal map guidance of ControlNet [102]. In Fig. 6 (b) , pre-trained text-to-image priors often produce renderings that deviate from the original identity. In Fig. 6 (c), personalized diffusion priors improve identity preservation but struggle with appearance consistency, as they lack input image information to hallucinate novel views. Our approach learns to jointly generate multiple novel views conditioned on the input image, thus achieving higher realism and view consistency in both identity and appearance.
What is the effect of normal map conditioning for multi-view diffusion models?
Our multi-view head diffusion models condition on normal maps of FLAME reconstruction, which are pixel-aligned with the target novel-view images. Fig. 6 (d) and (e) reflect that using camera pose embedding and ray map conditioning could introduce obvious misalignment errors in synthesizing pseudo-image ground truths, leading to blurred 3D Gaussian renderings.
@@
We can instead use Score Distillation Sampling (SDS) loss to constrain multi-view renderings. As shown in Fig. 6 (f), vanilla one-step SDS loss has appearance oversaturation issues in the face region.
@@
Through Fig. 6 (g) vs. (h), we can see that the upsampler module significantly sharpens and enhances facial appearance details.
5.3 Multi-view Head Diffusion Models
We compare normal map conditioning to alternatives in our multi-view head diffusion models, including pose embedding and Plücker ray maps [57]. Results in Tab. 4 and Fig. 7 show that FLAME normal maps can achieve more superior facial alignment, with ×$\times 2.56$ lower $l_{1}$ error and $7.73\%$ higher SSIM.
Method $l_{1}$downarrow LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$
pose embedding 0.123 0.184 16.52 76.12
ray map 0.137 0.198 16.05 75.71
normal map 0.048 0.119 19.67 83.85
Table 4: Ablation studies on different camera conditioning in multi-view head diffusion.
Figure 7: Ablation studies on different camera conditioning in multi-view head diffusion. We show predictions in the first row and $l_{1}$ difference maps between the ground truth and predictions.
@@ 6 Conclusion
In this work, we present a novel method to reconstruct photo-realistic head avatars from monocular videos to push the frontier of avatar fidelity from commodity devices. Due to limited observation and coverage of human heads, Gaussian reconstruction from monocular videos is inherently under-constrained. To address this challenge, we introduce multi-view diffusion priors that jointly constrain photorealism across multiple views rendered from Gaussian splats. We obtain these priors by designing a multi-view head diffusion model, fine-tuned on a multi-view head video dataset to generate novel views from a single image, conditioned on rendered normal maps from FLAME tracking. To mitigate over-saturated appearances, we use iteratively denoised images as pseudo-ground truths. A pre-trained latent upsampler diffusion model further enhances facial details. By reducing data capture requirements for avatar creation, our approach has the potential to unlock new opportunities in immersive VR/AR applications and products.
@@ Acknowledgements
This work was supported by Toyota Motor Europe and Woven by Toyota. This work was also supported by the ERC Starting Grant Scan2CAD (804724). We thank Rundi Wu, Simon Giebenhain, and Lei Li for constructive discussions and Angela Dai for the video voice-over.
@@ Appendix A Dataset
The statistics of the monocular video data are summarized in  Tab. 6. Since monocular videos captured on commodity devices lack corresponding ground truths for novel view renderings, we only evaluate avatar animation performance in the quantitative comparisons, by applying pose and expression parameters from those unseen frames during training.
@@ B.1 Monocular Head Tracking
The tracking is performed by minimizing a combination of multiple energy terms: (i) a photometric energy term, computed between the rendered image and the ground-truth one; (ii) a landmarks energy term, which computes the distance between the projected 2D FLAME [40] landmarks and the 2D landmarks predicted by an off-the-shelf detector [108]; (iii) temporal energy terms, applied on the per-timestep parameters, which ensure smoothness over time; (iv) regularization energy terms, applied on all FLAME [40] parameters. We revised the loss weights for the smoothness terms as: $\lambda_{smooth,transl}={3}\mathrm{e}{4}$, $\lambda_{smooth,rot}={3}\mathrm{e}{3}$, $\lambda_{smooth,jaw}=4.0$, $\lambda_{smooth,eyes}=1.0$, $\lambda_{smooth,expr}=0.5$. For all the remaining hyper-parameters we refer to the original work [1].
@@ B.2 Multi-view Latent Head Diffusion
Figure 9: Network architecture of our multi-view head latent diffusion model. The denoiser network is built on a 2D U-Net architecture with attention blocks. The input consists of multi-view image latents concatenated with VAE latents of normal maps rendered from the FLAME mesh. The 3D Attention block enforces 3D consistency by applying cross-attention across all views. It also incorporates the input image latent into the denoising process, effectively preserving the identity and appearance details of the input portrait.
B.3 Gaussian Regularizations
The position regularization term ensures that Gaussians remain close to their attached triangles during optimization through:
It will be disabled when the local scale of the Gaussian w.r.t the attached triangle is less than $\epsilon_{pos}=0.6$.
B.4 Ablation Studies of Gaussian Avatars Fusion
In the main paper, we evaluate various diffusion priors for novel view constraints, demonstrating the effectiveness of our face-specific multi-view diffusion priors for 3D Gaussian head reconstruction from monocular videos. Here, we provide additional details on the implementation of our ablation studies. We use six sequences from the NeRSemble dataset, including ’055 EXP-5’, ’098 EMO-1’, ’134 EMO-1’, ’165 EMO-1’, ’221 EXP-8’, and ’417 EMO-4’.
No diffusion.
This variant does not apply any priors to constrain novel view renderings. It is implemented using GA [60] with $SH=$0.
Pretrained Stable Diffusion.
This variant randomly renders a novel view at each iteration and refines it using Stable Diffusion 2.1, guided by ControlNet [102] with normal maps. The iteratively denoised images serve as pseudo-ground truths.
Personalized Stable Diffusion.
Instead of using a pre-trained model, this variant employs DreamBooth [66] to fine-tune the U-Net and text encoder with a learning rate of ${5}\mathrm{e}{-6}$ for 500 iterations. The iteratively denoised images are used as pseudo-supervision.
Pose-conditioned multi-view diffusion.
This variant uses pose embedding-conditioned multi-view diffusion models to generate pseudo-ground truths.
Raymap-conditioned multi-view diffusion.
This variant uses ray map-conditioned multi-view diffusion models to generate pseudo-ground truths.
Our multi-view diffusion using Score Distillation Sampling (SDS) loss.
Instead of using iteratively denoised images as pseudo-ground truths, this variant employs SDS loss [58] based on single-step denoising.
Ours without latent upsampler $\times$2.
We remove the pretrained latent upsampler $\times$2. Then the resolution of pseudo ground truths is 256$\times$256.
Ours without 3D-aware denoising.
The pseudo-ground truths generated by diffusion models rely on multi-view renderings of 3D Gaussian avatars, embedding 3D awareness into the diffusion process. To eliminate this 3D awareness, we set the time step to 0 when adding noise. Comparisons in Tab. 7 demonstrate the benefits of 3D-aware denoising.
Method Novel Views Novel Expressions
Ours, w/o 3D-aware 0.119 21.52 89.71 0.079 25.20 91.99
Ours final 0.118 21.82 89.87 0.079 25.39 92.02
Table 7: Ablation study of 3D-aware denoising.
Appendix C Results of Multi-view Head Diffusion
@@
Figure 10: Generation sample results of the multi-view head latent diffusion model. Given a single image as input, our method can generate identity-preserved, view-consistent multi-view portrait images.
@@ Appendix D Additional Comparisons
D.1 NeRSemble Dataset
In Fig. 11, we provide additional qualitative comparisons on dynamic head avatar reconstruction from monocular videos sampled from NerSemble dataset [36].
D.2 Monocular Videos on Commodity Devices
In Fig. 12, we provide additional qualitative comparisons against INSTA [110], FlashAvatar [92], and GA [60] on monocular videos captured by commodity devices.
Input GT PanoHead SphereHead INSTA FlashAvatar GA P4D-v2 GAGAvatar Ours
Figure 11: Additional results on novel view synthesis from monocular videos from the NeRSemble dataset. Our method demonstrates robust reconstruction of less observed regions (e.g., side facial areas), maintains facial identities across viewpoints, and consistently produces more plausible and view-consistent renderings from hold-out views.
(a) Input (b) GT (c) INSTA (d) FlashAvatar (e) GA (f) Ours
Figure 12: Additional comparisons of head avatar reconstruction from monocular videos on commodity devices. We present novel expression animation results using unseen frames from the monocular videos during training. Additionally, we display the fitting results for the input frame (top right) and novel view renderings of the same frame (bottom right). While all methods accurately fit observed frames in front-facing sequences with limited head poses, baseline methods fail to generalize to novel views and poses due to the absence of effective priors for less unobserved regions.
@@ D.2 Monocular Videos on Commodity Devices
D.3 Self- & Cross-Reenactment